OpenStudy (anonymous):

Give an example of a function, g(x), that has a local maximum at (-3; 3) and a local minimum at (3, -3).

OpenStudy (amistre64):

g'(3) = -3 eh...

OpenStudy (amistre64):

its a cubic at its basic

OpenStudy (amistre64):

centered at the the origin

OpenStudy (amistre64):

g'(3) = 0 and g'(-3) = 0

OpenStudy (amistre64):

x^3 derived is 3x^2 3(3)^2 = 0 27 = 0 27 - 27 = 0

OpenStudy (amistre64):

3x^2 - 27 = g' x^3 - 27x +C = function of g(x); just gotta calibrate for (3,-3)

OpenStudy (amistre64):

-3 = (3)^3 -27(3) + C -3 -17 + 27(3) = C C = 61 g(x) = x^2 -27x + 61 perhaps?

OpenStudy (amistre64):

g(x) = x^3 -27x + 61 that is lol

OpenStudy (anonymous):

Thanks a lot. I really appreciate it.

OpenStudy (amistre64):

it might be wrong; i havent checked it good wnough yet :)

OpenStudy (amistre64):

C = 51 maybe...

OpenStudy (anonymous):

Hmm, when I graph it it the maximum and minimum isn't (-3, 3) and (3, -3) let me try with C=51

OpenStudy (amistre64):

i see it too; :) gotta figure out where I looked at it wrong

OpenStudy (amistre64):

g' = 3x^2 - 27 = 0 at -3 and 3 so the derivative is good

OpenStudy (amistre64):

g(x) = x^3 - 27x + C if we integrate it...

OpenStudy (amistre64):

(x=-3,y=3) should fit it and (x=3,y=-3) should fit it, so we solve for C with these parameters

OpenStudy (amistre64):

-3 = (3)^3 - 27(3) + C and 3 = (-3)^3 - 27(-3) + C -3 - (3)^3 + 27(3) = C OR 3 - (-3)^3 + 27(-3) = C

OpenStudy (amistre64):

C = 51 or C = -51 which doesnt work out to well in teh end

OpenStudy (amistre64):

g(0) would also have to equal 0

OpenStudy (amistre64):

C = 0

OpenStudy (anonymous):

I just graphed it again and the minimum is at (3, -3) but not the Maximum

OpenStudy (anonymous):

the maximum is not at (3, -3)

OpenStudy (amistre64):

g(x) = x^3 - 27x ... thats it

OpenStudy (amistre64):

ack!!... that aint it either.. lol

OpenStudy (amistre64):

soooo close

OpenStudy (anonymous):

Haha, it's ok, thanks a lot for the help though. But I got a little help from someone but have no idea how hey got it, maybe you can help elaborate this. For this problem they started from this: (x+3)(x-3) g'(x)= X^2 -9 = 1/3x^3 -9 ???? Have no idea as to how they got here for this problem.

OpenStudy (amistre64):

i got some idea, but have to look at it some more

OpenStudy (amistre64):

g'(x) has roots at 3 and -3 so g'(x) = x^2 - 9 if we combine them; which is good; then integrate up to x^3 --- - 9x tho 3

OpenStudy (amistre64):

(1/3)(x^3) - 9x + C is the family of curves for it....

OpenStudy (amistre64):

C = -15 then i think (1/3)x^3 - 9x -15 ; lets check it

OpenStudy (anonymous):

Now the max is right but not the min...

OpenStudy (anonymous):

I did this very quickly; \(f(x)=x^3-27x-51\). Check it!!

OpenStudy (amistre64):

did that first lol

OpenStudy (amistre64):

it works for one , but not the other

OpenStudy (amistre64):

g(0) = 0 has to be part of te graph

OpenStudy (amistre64):

so adding somethig to it aint gonna help; we gotta find the right 'slope'

OpenStudy (amistre64):

M(x^3 - Bx) is what we need

OpenStudy (amistre64):

Dx(Mx^3 - MBx) 3Mx^2 - MB M(3x^2 - B) = 0 when x = 3 or -3

OpenStudy (amistre64):

Maybe include more powers of x?

OpenStudy (amistre64):

or derive again for inflection...

OpenStudy (amistre64):

6Mx = 0 when x = 0 lol ... that dint work out to well

OpenStudy (anonymous):

haha, thanks, i really appreciate the help

OpenStudy (amistre64):

k(-3)^3 = 3 -27k = 3 k = -1/9 maybe? x^3 ---- ; 3x^2/9 = x^2/3 i wonder 9